### v1

@@ 6 Symmetries of the mass matrices
where $F_{q}\simeq(m^{q}_{1}-m^{q}_{2})(m^{q}_{2}-m^{q}_{3})(m^{q}_{3}-m^{q}_{1})/(m_{3}^{q})^{3}$. The determinant ${\cal D}$ is readily shown to flip sign under either of the two $\bm{\lambda_{q}}$ sign reversals separately (given a $\bm{\lambda_{q}}$ phase difference of $\tfrac{\pi}{2}$), and of course under $CP$ transformation. Interestingly, this latter check only works in the case of a texture in which the 13 and 31 elements of the MMs are zero (or have cubic powers in their respective $\bm{\lambda_{q}}$), linking this invariance also to the texture zeroes. Thus, it seems that requiring invariance of the observables under the compound transformation also helps to constrain the form of the texture, beyond just the phase relationship of the $\bm{\lambda_{q}}$.
We turn now to consider how the ratio $\lambda_{u}/\lambda_{d}$ may be constrained to the value $\tan{\tfrac{\pi}{8}}$ by means of another invariance of the MMs. It is useful first to consider the more general situation where this is not the case (but the $\bm{\lambda_{q}}$ still retain their $\tfrac{\pi}{2}$ phase difference). Again the UT and Jarkskog’s determinant are instrumental to the argument. Figure 3 illustrates the LO-UT for this relaxed texture,
Figure 3: Leading-order UT, for our texture-pair (yellow, $\tilde{\beta}=\tfrac{\pi}{8}$), and (unshaded) three possible generalisations which represent a continuum with variable opening angle $\tilde{\beta}$ in the range $-\tfrac{\pi}{2}<\tilde{\beta}<\tfrac{\pi}{2}$ (negative values flip the sign of $J_{CP}$). The $CP$-conjugate LO-UT ($\tilde{\beta}=-\tfrac{\pi}{8}$) is shown in green. It is reachable by the particular $\tilde{\beta}$ rotation: $\tilde{\beta}\rightarrow\tilde{\beta}-\tfrac{\pi}{4}$, but only because $\tilde{\beta}=\beta_{0}=\tfrac{\pi}{8}$. The figure is drawn in the PDG phase convention, although, since $\tilde{\beta}$ is a relative phase, the figure can be rotated between phase conventions. Note that $|\bm{\overline{z}_{0}}+\bm{z_{0}}|=1$ for all $\tilde{\beta}$ and for all phase conventions.
and we see that:
@@
${\cal D}=-i\lambda_{0}^{10}b^{2}A_{0}^{2}\sin^{3}{2\tilde{\beta}}f_{{\cal D}}(b,c_{u},c_{d},\lambda_{0},\cos{2\tilde{\beta}})$ (25)
@@ 7 Discussion
In summary, we have introduced a novel texture for the two normalised quark mass matrices, having five free parameters in total. We use a small expansion parameter similar to Wolfenstein’s, but, motivated by the unitarity triangle, split in this case into two distinct complex components, $\bm{\lambda_{u}}$ and $\bm{\lambda_{d}}$, whose complex sum is $\lambda+{\cal O}(\lambda^{3})$. These complex components have ratio $\bm{\lambda_{u}}/\bm{\lambda_{d}}=-i\tan{\tfrac{\pi}{8}}$, which yields a UT consistent with experiment. In this “geomentric” ansatz, the distinctiveness of the mass hierarchies for the two quark sectors, that of the “up” quarks being more extreme than that of the “down” quarks, is directly linked to the different magnitudes of the two sloped sides of the UT (i.e. $|V_{ub}|<|V_{td}|$) via the relation $|\bm{\lambda_{u}}/\bm{\lambda_{d}}|\simeq|V_{ub}|/|V_{td}|\simeq\tan\beta<1$, which controls both. We obtain a good fit to the data renormalised to a scale in the range $(0.3-3)\times 10^{4}$ TeV at 68% CL.
We do not propose an extension of the SM with additional fields and quantum numbers to underpin the proposed texture, restricting our analysis to the phenomenology of the MMs themselves. Thus, we have instead elucidated two distinct new symmetries, manifested by the MMs, but absent from the SM, each involving an elementary transformation of the $\bm{\lambda_{q}}$ alone, with symmetry restored in each case by a $CP$ transformation. We have shown that these symmetries are necessary and sufficient to constrain the UT angles to the experimentally-determined values. Whether these symmetries occur in isolation, or are a manifestation of some higher symmetry is, for the time-being, unknown.
The up/down (weak isospin) reflection symmetry among the coefficients of the mass matrices is needed to reduce the number of parameters to make the ansatz more predictive, and is consistent with the data. The fact that two distinct parameters are needed in the smallest ‘11’ entry of the MMs is the least appealing aspect of the texture, but it is perhaps not entirely surprising, since radiative corrections due to the gauge interactions might dominate over the Yukawa contribution in this very small element and cannot be expected to be isospin-symmetric in the normalised mass matrices. Notwithstanding this feature, the proposed textures fit the experimental results and reduce the number of parameters needed to describe the data in the quark flavour sector from ten to seven (including the normalisations of the MMs).
@@ Acknowledgements
It is a pleasure to thank Tim Gershon, Rama Krishnan, Chris Sachrajda, Bill Scott and Jack Shergold for useful discussions and/or helpful comments on the manuscript. PFH acknowledges support from the UK Science and Technology Facilities Council (STFC) on the STFC consolidated grant ST/H00369X/1.
@@ Appendix A Analytic NLO solutions
$\lambda$ $=\lambda_{0}\left[1+\tfrac{f_{\lambda}}{4b}\lambda_{0}^{2}\,\right]+{\cal O}(\lambda_{0}^{5})$ (27)
$A$ $=A_{0}\left[1+\tfrac{1}{4b}(3b^{2}-2b\rho_{0}-2f_{\lambda})\lambda_{0}^{2}\,\right]+{\cal O}(\lambda_{0}^{4})$ (28)
$\overline{\rho}$ $=\rho_{0}\left\{1-\tfrac{1}{8b\rho_{0}}\left[f_{b}+2\eta_{0}(2f_{a}-2b-7b^{2})\right]\lambda_{0}^{2}\,\right\}+{\cal O}(\lambda_{0}^{4})$ (29)
$\overline{\eta}$ $=\eta_{0}\left\{1-\tfrac{1}{4b}\left[2f_{a}-4b+3b^{2}+4\eta_{0}f_{b}\right]\lambda_{0}^{2}\,\right\}+{\cal O}(\lambda_{0}^{4})$ (30)
where $f_{\lambda}=(3f_{a}-5b+4\eta_{0}\,\delta_{c})$, $f_{a}=A_{0}^{2}+\tfrac{1}{2}(c_{d}+c_{u})$, $f_{b}=b+\delta_{c}$ and $\delta_{c}=c_{d}-c_{u}$. The NLO corrections in eqs. (27)-(30) as fractions of the LO terms (calculated using the fitted values reported in table 1) are respectively: $-5.8\times 10^{-3}$, $+2.6\%$, $+3.6\%$ and $-1.8\%$. For the mass ratios, we find:
@@
where $r_{q}=\tfrac{c_{q}}{b}$ and $r_{A}=\tfrac{A_{0}^{2}}{b}$. The NLO corrections to the mass ratios $m_{c}/m_{t}$, $m_{s}/m_{b}$, $m_{u}/m_{c}$, $m_{d}/m_{s}$ as a fraction of the LO terms are respectively $-4.3\times 10^{-3}$, $-2.5\%$, $+4.3\%$, and $+4.5\%$ (calculated using the fitted values reported in table 1). In all cases, the results are compatible with the full numerical results reported in table 1, given the indicated degree of approximation.

### v4

@@ 6 Symmetries of the mass matrices
where $F_{q}\simeq(m^{q}_{1}-m^{q}_{2})(m^{q}_{2}-m^{q}_{3})(m^{q}_{3}-m^{q}_{1})/(m_{3}^{q})^{3}$. The determinant ${\cal D}$ is readily shown to flip sign under either of the two $\bm{\lambda_{q}}$ (i.e. $\bm{z_{0}}$ or $\bm{\overline{z}_{0}}$) sign reversals separately (given a $\bm{\lambda_{q}}$ phase difference of $\tfrac{\pi}{2}$), and of course under $CP$ transformation. Interestingly, this latter check only works in the case of a texture in which the 13 and 31 elements of the MMs are zero, linking this invariance also to the texture zeroes. Thus, it seems that requiring invariance of the observables under the compound transformation also helps to constrain the form of the texture, beyond just the phase relationship of the $\bm{\lambda_{q}}$.
Figure 3: Leading-order UT, for our texture-pair (yellow, $\tilde{\beta}=\tfrac{\pi}{8}$), and (unshaded) three possible generalisations which represent a continuum with variable opening angle $\tilde{\beta}$ in the range $-\tfrac{\pi}{2}<\tilde{\beta}<\tfrac{\pi}{2}$ (negative values flip the sign of $J_{CP}$). The $CP$-conjugate LO-UT ($\tilde{\beta}=-\tfrac{\pi}{8}$), in green, is reachable by the particular $\tilde{\beta}$ transformation, eq. (26), i.e. $\tilde{\beta}\rightarrow\tilde{\beta}-\tfrac{\pi}{4}$, but only because $\tilde{\beta}=\beta_{0}=\tfrac{\pi}{8}$. The figure is drawn in the PDG phase convention, although, since $\tilde{\beta}$ is a relative phase, the figure can be rotated between phase conventions. Note that $|\bm{\overline{z}_{0}}+\bm{z_{0}}|=1$ for all $\tilde{\beta}$ and for all phase conventions.
We turn now to consider how the ratio $\lambda_{u}/\lambda_{d}$ may be constrained to the value $\tan{\tfrac{\pi}{8}}$ by means of another invariance of the MMs. It is useful first to consider the more general situation where this is not the case (but the $\bm{\lambda_{q}}$ still retain their $\tfrac{\pi}{2}$ phase difference). Again the UT and Jarkskog’s determinant are instrumental to the argument. Figure 3 illustrates the LO-UT for this relaxed texture, and we see that:
@@
${\cal D}=-i\lambda_{0}^{10}b^{2}A_{0}^{2}\sin^{3}\!{(2\tilde{\beta})}f_{{\cal D}}(b,c_{u},c_{d},\lambda_{0},\cos{2\tilde{\beta}})$ (25)
@@ 7 Discussion
In summary, we have introduced a novel texture for the two normalised quark mass matrices, having five free parameters in total. We use a small expansion parameter similar to Wolfenstein’s, but, motivated by the UT, split here into two distinct complex components, $\bm{\lambda_{u}}$ and $\bm{\lambda_{d}}$, whose complex sum has magnitude $\lambda+{\cal O}(\lambda^{3})$. These complex components have a fixed ratio $\bm{\lambda_{u}}/\bm{\lambda_{d}}=-i\tan{\tfrac{\pi}{8}}$, which yields a UT consistent with experiment. In this “geometric” ansatz, the distinctiveness of the mass hierarchies for the two quark sectors, that of the “up” quarks being more extreme than that of the “down” quarks, is directly linked to the different magnitudes of the two sloped sides of the UT (i.e. $|V_{ub}|<|V_{td}|$) via the relation $|\bm{\lambda_{u}}/\bm{\lambda_{d}}|\simeq|V_{ub}|/|V_{td}|\simeq\tan\beta<1$, which controls both. We obtain a good fit to the data renormalised to a scale in the range $(0.3-3)\times 10^{4}$ TeV.
We do not propose an extension of the SM with additional fields and quantum numbers to underpin the proposed texture, restricting our analysis to the phenomenology of the MMs themselves. We have instead elucidated two distinct new symmetries, manifested by the MMs, but absent from the SM, each involving an elementary transformation of the $\bm{\lambda_{q}}$ alone, with symmetry restored in each case by a $CP$ transformation. We have shown that these symmetries are necessary and sufficient to constrain the UT angles to the experimentally-determined values. Whether these symmetries occur in isolation, or are a manifestation of some higher symmetry is, for the time being, unknown.
The weak isospin reflection symmetry among the coefficients of the MMs is needed to reduce the number of parameters to make the ansatz more predictive, and is consistent with the data. The fact that two distinct parameters are needed in the smallest ‘11’ entry of the MMs is the least appealing aspect of the texture, but it is perhaps not entirely surprising, since radiative corrections due to the gauge interactions might dominate the Yukawa contribution in this very small element and cannot be expected to be isospin-symmetric in the normalised mass matrices. Notwithstanding this feature, the proposed textures fit the experimental results and reduce the number of parameters needed to describe the data in the quark flavour sector from ten to seven (including the normalisations of the MMs).
@@ Acknowledgements
It is a pleasure to thank Tim Gershon, Rama Krishnan, Chris Sachrajda and Jack Shergold for useful discussions and/or helpful comments on the manuscript. PFH acknowledges support from the UK Science and Technology Facilities Council (STFC) on the STFC consolidated grant ST/W000571/1.
@@ Appendix A Analytic NLO solutions
$\lambda$ $=\lambda_{0}\left(1+f_{\lambda}\lambda_{0}^{2}\,\right)+{\cal O}(\lambda_{0}^{5})$ (27)
$A$ $=A_{0}\left\{1+\left[\tfrac{1}{4}(3b-2\rho_{0})-2f_{\lambda}\right]\lambda_{0}^{2}\,\right\}+{\cal O}(\lambda_{0}^{4})$ (28)
$\overline{\rho}$ $=\rho_{0}\left(1+c_{0}f_{\rho}\lambda_{0}^{2}\,\right)+{\cal O}(\lambda_{0}^{4})$ (29)
$\overline{\eta}$ $=\eta_{0}\left\{1+\left[s_{0}f_{\rho}+\tfrac{1}{2}(1-5b)\right]\lambda_{0}^{2}\,\right\}+{\cal O}(\lambda_{0}^{4}),$ (30)
where $f_{\lambda}\equiv\tfrac{1}{4}(3f_{A}+4\eta_{0}f_{c}-5)\simeq-0.11$, $f_{A}\equiv\tfrac{1}{b}\!\left[A_{0}^{2}+\tfrac{1}{2}(c_{d}+c_{u})\right]$, $f_{c}\equiv\tfrac{1}{b}(c_{d}-c_{u})$, and $f_{\rho}\equiv s_{0}(1+f_{c})-\tfrac{1}{2s_{0}}\!(f_{A}+f_{c}-\tfrac{7}{2}b)\simeq 0.76$. For the mass ratios, we find:
@@
where $r_{q}=\tfrac{c_{q}}{b}$ and $r_{A}=\tfrac{A_{0}^{2}}{b}$. The NLO corrections in eqs. (27)-(30) as fractions of the LO terms are respectively: $-5.8\times 10^{-3}$, $+2.6\%$, $+3.6\%$ and $-1.8\%$, while the NLO corrections in eqs. (31)-(32) to the mass ratios $m_{c}/m_{t}$, $m_{s}/m_{b}$, $m_{u}/m_{c}$, $m_{d}/m_{s}$ as a fraction of the LO terms are respectively $-4.3\times 10^{-3}$, $-2.5\%$, $+4.3\%$, and $+4.5\%$ (all calculated using the fitted values reported in table 1). In all cases, the results are compatible with the full numerical results reported in table 1, given the indicated degree of approximation.